OpenStudy (anonymous):

How do I make a table of solutions to graph for 2x + y = 3

OpenStudy (saifoo.khan):

When x-intercepts, make y=0 When y-intercepts, make x = 0

OpenStudy (anonymous):

What?

OpenStudy (saifoo.khan):

When you have to find x-intercepts, make y=0 2x+y = 3 2x +0 = 3 2x = 3 x =3/2 ------------ Same with y-intercept. When y-intercepts, make x = 0

OpenStudy (anonymous):

How do I find more solutions though. If I do only that all I'll be getting is 3/2.